# Retrieval-Augmented Generation

### v1

@@ 1 Introduction
Our design of MiniRAG is motivated by three fundamental observations: (1) while SLMs struggle with semantic understanding, they excel at pattern matching and localized text processing; (2) explicit structural information can compensate for limited semantic capabilities by providing navigational cues for retrieval; and (3) decomposing complex RAG operations into simpler, well-defined steps can maintain system robustness without requiring advanced reasoning capabilities. These insights lead us to prioritize structural knowledge representation over semantic complexity, and to leverage graph-based patterns that naturally align with SLMs’ strengths while circumventing their limitations. This design philosophy enables MiniRAGto achieve efficient and reliable performance even with lightweight models, marking a significant departure from traditional LLM-centric RAG architectures.
@@ 2.1 Heterogeneous Graph Indexing with Small Language Models
In resource-constrained RAG systems, Small Language Models (SLMs) introduce significant operational constraints that impact their effectiveness. These limitations primarily manifest in two critical areas: i) reduced capability to extract and understand complex entity relationships and subtle contextual connections from raw text, and ii) diminished capacity to effectively summarize large volumes of text and process retrieved information containing noise and irrelevant content.
@@ 2.2.1 Query Semantic Mapping
In the retrieval phase, the primary goal for a user-input query $q$ is to identify elements related to the query (e.g., text chunks) from the constructed index data, thereby aiding the model in generating accurate responses. To achieve this, it is essential to first parse the query and align it with the index data. Some prior RAG methods utilize LLMs to expand or decompose the query into fine-grained queries (Chan et al., 2024; Edge et al., 2024a; Guo et al., 2024), enhancing the match between the query and the index data. However, this process relies on LLMs to extract high-quality abstract information from the query, which poses challenges for smaller language models. Therefore, in the retrieval process of MiniRAG, we leverage entity extraction—a relatively simple and effective task for small language models—to facilitate the decomposition and mapping of the query $q$ to our graph-based indexing data (i.e., the semantic-aware heterogeneous graph $\mathcal{G}$).
@@
Query-Driven Reasoning Path Discovery. Within our semantic-aware heterogeneous graph $\mathcal{G}={\mathcal{V}_{c},\mathcal{V}_{e}}$, $\mathcal{M}$ constructs reasoning paths through an intelligent query-guided mechanism. For any input query $q$, the model identifies relevant text chunks by jointly considering two key aspects: (1) semantic relevance between query and entity nodes, and (2) structural coherence among entity-entity and entity-chunk relationships. This dual-objective optimization framework simultaneously maximizes $q$-$\mathcal{V}{e}$ semantic alignment and preserves ($\mathcal{V}_{e}$-$\mathcal{V}_{e}$), ($\mathcal{V}_{e}$-$\mathcal{V}_{c}$) relational dependencies, effectively capturing complex reasoning chains within the heterogeneous knowledge graph. The systematic query-relevant reasoning path discovery procedure consists of the following key steps:
@@ 2.2.2 Topology-Enhanced Graph Retrieval
  • Key Relationship Identification: We first identify high-quality entity-entity connections within graph $\mathcal{G}$ that are relevant to query $q$ through node-edge interactions. In the entity-entity connections $\mathcal{E}_{\alpha}$, we define an edge as highly relevant if it connects a starting node $\hat{v}_{s}\in\hat{\mathcal{V}}_{s}$ to an answer node $\hat{v}a\in\hat{\mathcal{V}}a$ along their shortest path. For efficient extraction, we focus on edges proximate to starting or answer nodes. The relevance scoring function $\omega_{e}(\cdot)$ for each edge $e\alpha\in\mathcal{E}{\alpha}$ is formally defineds:
@@ 3.2 Performance Analysis (RQ1)
$\bullet$ Storage Efficiency While Maintaining Performance. MiniRAG demonstrates exceptional storage efficiency while preserving high accuracy levels. Empirical evaluations show that MiniRAG achieves superior accuracy while requiring only 25% of the storage space compared to baselines like LightRAG with gpt-4o-mini. This dramatic reduction in storage requirements, coupled with maintained or improved accuracy, represents a significant advancement in efficient RAG system design.

### v3

@@ 2.1 Heterogeneous Graph Indexing with Small Language Models
In resource-constrained RAG systems, Small Language Models (SLMs) introduce significant operational constraints that impact their effectiveness. These limitations primarily manifest in two critical areas: i) reduced capability to extract and understand complex entity relationships and subtle contextual connections from raw text corpus, and ii) diminished capacity to effectively summarize large volumes of text and process retrieved information containing noise and irrelevant content.
@@ 2.2.1 Query Semantic Mapping
In the retrieval phase, the primary goal for a user-input query $q$ is to identify elements related to the query (e.g., named entities, text chunks) from the constructed index data, thereby aiding the model in generating accurate responses. To achieve this, it is essential to first parse the query and align it with the index data. Some prior RAG methods utilize LLMs to expand or decompose the query into fine-grained queries (Chan et al., 2024; Edge et al., 2024a; Guo et al., 2024), enhancing the match between the query and the index data. However, this process relies on LLMs to extract high-quality abstract information from the query, which poses challenges for smaller language models. Therefore, in the retrieval process of MiniRAG, we leverage entity extraction—a relatively simple and effective task for small language models—to facilitate the decomposition and mapping of the query $q$ to our graph-based indexing data (i.e., the semantic-aware heterogeneous graph $\mathcal{G}$).
@@
Query-Driven Reasoning Path Discovery. Within our semantic-aware heterogeneous graph $\mathcal{G}=\{\mathcal{V}_{c},\mathcal{V}_{e}\}$, MiniRAG constructs reasoning paths through an intelligent query-guided mechanism. For any input query $q$, the model identifies relevant text chunks by jointly considering two key aspects: (1) semantic relevance between query and entity nodes, and (2) structural coherence among entity-entity and entity-chunk relationships. This dual-objective optimization framework simultaneously maximizes $q$-$\mathcal{V}{e}$ semantic alignment and preserves ($\mathcal{V}_{e}$-$\mathcal{V}_{e}$), ($\mathcal{V}_{e}$-$\mathcal{V}_{c}$) relational dependencies, effectively capturing complex reasoning chains within the heterogeneous knowledge graph. The systematic query-relevant reasoning path discovery procedure consists of the following key steps:
@@ 2.2.2 Topology-Enhanced Graph Retrieval
  • Key Relationship Identification: We first identify high-quality entity-entity connections within graph $\mathcal{G}$ that are relevant to query $q$ through node-edge interactions. In the entity-entity connections $\mathcal{E}_{\alpha}$, we define an edge as highly relevant if it connects a starting node $\hat{v}_{s}\in\hat{\mathcal{V}}_{s}$ to an answer node $\hat{v}_{a}\in\hat{\mathcal{V}}_{a}$ along their shortest path. For efficient extraction, we focus on edges proximate to starting or answer nodes. The relevance scoring function $\omega_{e}(\cdot)$ for each edge $e_{\alpha}\in\mathcal{E}_{\alpha}$ is formally defineds:
@@ 3.2 Performance Analysis (RQ1)
$\bullet$ Storage Efficiency While Maintaining Performance. MiniRAG demonstrates exceptional storage efficiency while preserving high accuracy levels. Empirical evaluations show that MiniRAG achieves competitive accuracy while requiring only 25% of the storage space compared to baselines like LightRAG w/ gpt-4o-mini. This dramatic reduction in storage requirements, coupled with maintained or improved accuracy, represents a significant advancement in efficient RAG system design.